solve the initial value problem:

dx/dt+ (tant)x= cos2t , x(0)= -1


1
Expert's answer
2021-07-11T17:58:23-0400

Integrating factor: μ(t)=sect\mu(t)=\sec t


sectdxdt+sect(tant)x=cost\sec t\dfrac{dx}{dt}+\sec t(\tan t)x=\cos t

ddt(xsect)=cost\dfrac{d}{dt}(x\sec t)=\cos t

xsect=costdtx\sec t=\int\cos t dt

xsect=sint+Cx\sec t=\sin t+C

x(0)=1x(0)=-1


1sec(0)=sin(0)+C-1\cdot\sec(0)=\sin (0)+C


C=1C=-1

xsect=sint1x\sec t=\sin t-1

Or


x=sintcostcostx=\sin t\cos t-\cos t